Whats the vertex form equation of this parabola?
Vertex: (7,-6)
Focus: (57/8,-6)

Answer by lwsshak3(11628)

Whats the vertex form equation of this parabola?
Vertex: (7,-6)
Focus: (57/8,-6)
***
Parabola opens rightward:
Its vertex form of equation: x=A(y-k)^2+h, (h,k)=coordinates of the vertex, A is a coefficient that affects the slope or steepness of the curve.
Basic form of equation:
(y-k)^2=4p(x-h)
(y+6)^2=4p(x-7)
p=1/8 (distance from vertex to focus on the axis of symmetry: (57/8)-7=1/8
4p=1/2
basic form of equation:(y+6)^2=(x-7)/2
vertex form of equation: x=2(y+6)^2+7